Circular No. 9007

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 Y1 (BOATTINI) A. Boattini reports his discovery of moderately-condensed comet of size

10′′ and with a tail ∼ 20′′ long in p.a. 70^{o} on images taken with the Catalina Sky Survey 0.68-m Schmidt telescope (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other astrometric CCD observers have also noted the object’s cometary appearance. F. Hormuth (Calar Alto, Spain, 1.23-m reflector) reports a 4′′ coma with a 15′′ tail in p.a. ∼ 80^{o} in five stacked 75-s unfiltered exposures from Dec. 22.76–22.77 UT; a stacked 15-min exposure from 24 hr later reveals a coma diameter of 8′′ and a tail 30′′ long in p.a. ∼ 70^{o}. R. Ligustri (Talmassons, Udine, Italy, 0.35-m f/5 reflector) reports that twenty 120-s exposures on Dec. 23.7 shows an 18′′ tail in p.a. 51^{o}. G. Masi (Ceccano, Italy, Dec. 24.74) finds the object to be diffuse with an extension extending to ∼ 5′′. L. Buzzi (Varese, Italy, 0.60-m f/4.6 reflector, Dec. 24.74–24.76) notes that stacked exposures show a compact coma 10′′ wide with a fanlike tail 15′′–20′′ long centered at p.a. ∼ 55^{o}–60^{o}, adding that a knot or condensation in the tail appears 10′′ from the nuclear condensation in p.a. 52^{o}; 25-min stacked exposures from Dec. 25.8 show a 10′′ coma elongated for 15′′ in p.a. ∼ 65^{o}, again suggesting that other knots or condensations appear to be following the main component. G. Hug (Scranton, KS, U.S.A., 0.56-m reflector, Dec. 25.1) finds the object to be somewhat diffuse.

2008 UT  α_{2000}  δ_{2000}  Mag.
Dec. 22.08726  22^{h}13^{m}51.^{s}56  −2^{o}54′17′′.3  17.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-Y53.

T = 2009 Feb. 16.808 TT  ω = 154.^{o}819
Ω = 255.500 }2000.0
q = 1.40724 AU  i = 8.959

COMET P/2008 X4 (CHRISTENSEN) J. N. Marcus, St. Louis, MO, U.S.A., suggested last week that comet

P/2008 X4 would be a good candidate for enhanced brightness due to significant forward-scattering geometry as its solar elongation approaches zero now (minimum elongation = 0^{o}.4 on Dec. 23.64 TT) — perhaps reaching total mag ∼ 4 — and he and K. Battams (NRL) note that the comet is visible clearly in SOHO/LASCO data from the C3 and C2 coronographs.

2008 December 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green